OpenStudy (anonymous):

Will give medals Is x - 8 a factor of the function f(x) = -2x3 + 17x2 - 64? Explain.

OpenStudy (anonymous):

these are the answers Yes. When the function f(x) = -2x3 + 17x2 - 64 is divided by x - 8, the remainder is zero. Therefore, x - 8 is a factor of f(x) = -2x3 + 17x2 - 64. No. When the function f(x) = -2x3 + 17x2 - 64 is divided by x - 8, the remainder is zero. Therefore, x - 8 is not a factor of f(x) = -2x3 + 17x2 - 64. Yes. When the function f(x) = -2x3 + 17x2 - 64 is divided by x - 8, the remainder is not zero. Therefore, x - 8 is a factor of f(x) = -2x3 + 17x2 - 64. No. When the function f(x) = -2x3 + 17x2 - 64 is divided by x - 8, the remainder is not zero. Therefore, x - 8 is not a factor of f(x) = -2x3 + 17x2 - 64.

myininaya (myininaya):

x-8=0 => x=8 plug in x=8 and see if you get 0 back

myininaya (myininaya):

if you get 0 back then on of the factors are x-8

OpenStudy (anonymous):

Put x=8 in the function or in other words evaluate f(8) if it equates to zero then it is a solution

myininaya (myininaya):

if not then one of the factors is not x-8

OpenStudy (anonymous):

so whats the answer ?

myininaya (myininaya):

tell me what you get for f(8)

myininaya (myininaya):

\[f(8)=-2(8)^3+17(8)^2-64=?\]

OpenStudy (anonymous):

Not not a factor

OpenStudy (anonymous):

i think its d

myininaya (myininaya):

you actually don't have to put this in a calculator to check \[f(8)=-2(8)^3+17(8)^2-8^2=8^2(-2(8)+17-1)=8^2(-16+17-1)\]

myininaya (myininaya):

is -16+17-1 zero?

OpenStudy (anonymous):

yes

myininaya (myininaya):

the answer is then x-8 is a factor of the function